OpenStudy (anonymous):

LAST Question of the day! how to convert from slope intercept form to standard form!

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

@Hero @JakeV8

OpenStudy (anonymous):

the equaion is y=3x+11

OpenStudy (aripotta):

do you know what standard form is?

OpenStudy (anonymous):

I think the answer is 3y-3x=11 but im not sure

hero (hero):

y = 3x + 11 -3x + y = 11 3x - y = -11

OpenStudy (anonymous):

yes its Ax+By=c

OpenStudy (aripotta):

well there you go...hero just gave you the answer :/

OpenStudy (anonymous):

Thank you!

hero (hero):

I gave him the steps to get to the answer

OpenStudy (anonymous):

her

hero (hero):

OOps

OpenStudy (anonymous):

wait so theres more?

hero (hero):

Nope, that's it

OpenStudy (anonymous):

oh okay lol